# P3-SAM: Native 3D Part Segmentation

### v1

@@ 1 Introduction
In this paper, we propose a native 3D Point-Promptable Part segmentation model termed P3-SAM, designed to fully automate the segmentation of any complex 3D objects into components with precise mask and strong robustness. As a pioneering promptable image segmentation work, SAM provides a feasible implementation approach.
However, our method focuses on achieving precise part segmentation automatically, and we simplify the architecture of SAM. Without adopting the complex segmentation decoder and multiple types of prompts from SAM, our model is designed to handle only one positive point prompt. Specifically, P3-SAM contains a feature extractor, three segmentation heads, and an IoU prediction head. We employ PointTransformerV3 Wu et al. (2024) as our feature extractor and integrate its features from different levels as extracted point-wise features. The input point prompt and feature are fused and passed to the segmentation heads to predict three multi-scale masks and an IoU (Intersection of Union) predictor is utilized to evaluate the quality of the masks. To automatically segment an object, we apply our segmentation model using point prompts sampled by FPS (Farthest Point Sampling) and utilize NMS (Non-Maximum Suppression) to merge redundant masks. The point-level masks are then projected onto mesh faces to obtain the part segmentation results.
@@ 3.1 Data Curation
Figure 2: Training pipeline of P3-SAM. Input point clouds are fed to feature extractor to obtain point features. The features, point prompts, and original point clouds are then fed to a two stage multi-mask segmentor to obtain the masks for the prompts. Finally, the IoU (Intersection of Union) predictor is utilized to evaluate the quality of the masks.
@@ 3.2.1 Network Architecture
Figure 3: Automatic segmentation pipeline. Point prompts are sampled by FPS (Farthest Point Sampling) and go through the P3-SAM to obtain multiple masks. NMS (Non-Maximum Suppression) is then adopted to merge redundant masks. The point-level masks are then projected onto mesh faces to obtain the part segmentation results.
@@ 4.1 Implementation Details
The channel of the point-wise feature $\mathbf{f}_{p}$ is $512$. The point number $N_{p}$ is set to $100,000$ during training, evaluation, and inference. We randomly select $K=8$ parts in the training process and set $\alpha_{dice}$ to $0.5$. For automatic segmentation, we sample $N_{pp}=400$ prompts from points, and the threshold $T_{NMS}$ is set to $0.9$. Our network is trained on our dataset using 64 H20 for 9 epochs. We set the batch size to 2 per GPU, and the training took approximately 4 days. We employ the Adam optimizer with a learning rate of $10^{-5}$.
@@ 4.2 Comparison
Table 1: The comparison of our method with related works on different aspects.
Methods Data Num. Data Type Param. Time(Seg.) Time(Inter.) Auto.
[…]
@@
The comparisons on PartObj-Tiny-WT and PartNetE are shown in Table 3. Since watertight data and point clouds lack connectivity information, PartField’s performance is not as good as on non-watertight data. This again validates that our method effectively learns the geometric features of objects. Another observation is that the metrics for interactive segmentation are lower than those for full segmentation. This is because, in the evaluation of interactive segmentation, the method can only rely on a single prompt point, focusing more on the accuracy of individual mask segmentation. This further validates the precision of our approach. Comparisons across various datasets and tasks, involving different data forms such as non-watertight meshes, watertight meshes, and point clouds, demonstrate the remarkable effectiveness, robustness and generalization ability of our methods, confirming its superior performance under diverse conditions. We also conduct a qualitative comparison of our method and previous methods, as shown in Figure 4. The results show that our method can predict masks with a reasonable number of categories, clear and accurate boundaries, and can handle complex models as well as different types of data and tasks. This demonstrates the great generalization and robustness of our approach.
Figure 4: The comparison of our method with previous methods on PartObj-Tiny and PartObj-Tiny-WT across different tasks including full segementation and interactive segmentation. Here, the green points indicate the input prompt pints.
@@ 4.3 Applications
Multi-Prompts Auto-Segmentation. Our method can also segment the object given several point prompts that indicate specific parts as shown in Figure 3. In this setting, we have a strong condition where each prompt corresponds to a specific part of the object. Therefore, instead of using the predicted IoU to evaluate mask quality, we only need to select one mask for each part such that all masks collectively cover the entire object as much as possible while minimizing overlap between the masks.
Hierarchical Part Segmentation. After segmentation, we can obtain the average feature of each part by finding the corresponding point-wise feature $\mathbf{f}_{p}$. We then directly employ hierarchical clustering on these average features to obtain hierarchical segmentation results. As shown in Figure 4, our hierarchical segmentation results effectively aggregate different parts at various levels, validating the effectiveness of our feature extraction method in accurately representing the information of each part.
@@
4.4 Ablation Study
We first conduct ablation studies on our network architecture. The feature extractor can be any point cloud encoder, and we select the current state-of-the-art one that is Sonata. The IoU predictor is critical for selecting the best mask; without it, our method cannot function properly. We then focus on conducting ablation studies of our two-stage multi-head segmentor. As shown in Table 4, we evaluate four ablated models and our full method on the test set of our dataset. The first model contains only one segmentation head of the first stage. The second and third models respectively include only the first and second stages. The fourth model includes both stages but is trained without data augmentation. The last model is our full version. This progressive ablation study clearly demonstrates the importance of each component in our method. Notably, the difference between the second and third models is that the latter extracts a global feature during segmentation. The better performance metrics of the third model highlight the importance of this global feature. As shown in Figure 5, we also present the full segmentation results without using the NMS or flood fill algorithm in our automatic segmentation approach. The results highlight the necessity of these two steps, as without them, the segmentation masks are not complete, have unclear boundaries, and do not yield a reasonable number of parts.

### v2

@@ 1 Introduction
In this paper, we propose a native 3D Point-Promptable Part segmentation model termed P3-SAM, designed to fully automate the segmentation of any complex 3D objects into components with precise mask and strong robustness. The improvements and differences of our method compared to previous methods are summarized in Table1. As a pioneering promptable image segmentation work, SAM provides a feasible implementation approach. However, our method focuses on achieving precise part segmentation automatically, and we simplify the architecture of SAM. Without adopting the complex segmentation decoder and multiple types of prompts from SAM, our model is designed to handle only one positive point prompt. Specifically, P3-SAM contains a feature extractor, three segmentation heads, and an IoU prediction head. We employ PointTransformerV3 Wu et al. (2024) as our feature extractor and integrate its features from different levels as extracted point-wise features. The input point prompt and feature are fused and passed to the segmentation heads to predict three multi-scale masks and an IoU (Intersection of Union) predictor is utilized to evaluate the quality of the masks. To automatically segment an object, we apply our segmentation model using point prompts sampled by FPS (Farthest Point Sampling) and utilize NMS (Non-Maximum Suppression) to merge redundant masks. The point-level masks are then projected onto mesh faces to obtain the part segmentation results.
@@
Table 1: The comparison of our method with related works across several key aspects, including the number and type of training data, the number of parameters, the time cost for full and interactive segmentation, and the ability to automatically segment objects.
Methods Data Num. Data Type Param. Time(Seg.) Time(Inter.) Auto.
[…]
@@ 3.1 Data Curation
Figure 2: The Network Architecture of P3-SAM. Input point clouds are fed to feature extractor to obtain point features. The features, point prompts, and original point clouds are then fed to a two stage multi-mask segmentor to obtain three masks in various scales for the prompts. Finally, the IoU predictor is utilized to evaluate the quality of the masks and select the best one as the final prediction.
@@ 3.3 Automatic Segmentation
Figure 3: Automatic Segmentation Pipeline. Point prompts are sampled by FPS and go through the P3-SAM to obtain multiple masks. NMS is then adopted to merge redundant masks. The point-level masks are then projected onto mesh faces to obtain the part segmentation results.
@@ 4.1 Implementation Details
The channel of the point-wise feature $\mathbf{f}_{p}$ is 512512. The point number $N_{p}$ is set to $100,000$ during training, evaluation, and inference. We randomly select $K=8$ parts in the training process and set $\alpha_{dice}$ to $0.5$. For automatic segmentation, we sample $N_{pp}=400$ prompts from points, and the threshold $T_{NMS}$ is set to $0.9$. Our network is trained on our dataset using 64 H20 for 9 epochs. We set the batch size to 2 per GPU, and the training took approximately 4 days. We employ the Adam optimizer with a learning rate of $10^{-5}$.
@@ 4.2 Comparison
Figure 4: The comparison of our automatic segmentation approach with previous methods on PartObj-Tiny.
@@
The comparisons on PartObj-Tiny-WT and PartNetE are shown in Table 3. Since watertight data and point clouds lack connectivity information, PartField’s performance is not as good as on non-watertight data. This again validates that our method effectively learns the geometric features of objects. Here, SAMesh will get stuck when processing watertight meshes due to the high number of faces. Another observation is that the metrics for interactive segmentation are lower than those for full segmentation. This is because, in the evaluation of interactive segmentation, the method can only rely on a single prompt point, focusing more on the accuracy of individual mask segmentation. This further validates the precision of our approach. Comparisons across various datasets and tasks, involving different data forms such as non-watertight meshes, watertight meshes, and point clouds, demonstrate the remarkable effectiveness, robustness and generalization ability of our methods, confirming its superior performance under diverse conditions.
We also conduct a qualitative comparison of our method and previous methods on PartObj-Tiny for full segmentation with connectivity, as shown in Figure 4. SAMesh tends to over-segment objects, while other methods struggle with handling complex objects, resulting in inaccurate masks, failure to separate multiple part masks, and other segmentation errors. However, our method can accurately segment complex objects, even performing part segmentation on the lizard and beetle in the scene of the last row. Figure 5 shows a comparison between our method and PartField on the PartObj-Tiny-WT dataset for full segmentation without connectivity. PartField struggles to segment watertight meshes, while our method can maintain the segmentation quality. Figure 5 also illustrates the interactive segmentation results of our method and Point-SAM given the green point prompts, where our segmentation results exhibit accurate boundaries and scales. The results show that our method can predict masks with a reasonable number of categories, clear and accurate boundaries, and can handle complex models as well as different types of data and tasks. They also demonstrate the great generalization and robustness of our approach.
Figure 5: The comparison of our method with previous methods on PartObj-Tiny and PartObj-Tiny-WT across different tasks including full segmentation and interactive segmentation. Here, the green points indicate the input prompt pints.
@@ 4.3 Applications
Multi-Prompts Auto-Segmentation. Our method can also segment the object given several point prompts that indicate specific parts. In this setting, we have a strong condition where each prompt corresponds to a specific part of the object. Therefore, instead of using the predicted IoU to evaluate mask quality, we only need to select one mask for each part such that all masks collectively cover the entire object as much as possible while minimizing overlap between the masks. As shown in Figure 6, the multi-prompts segmentation can follow the user’s instructions. Compared to automatic segmentation, it can both extend unsegmented regions, such as the horse’s body, and merge over-segmented regions, such as flower petals.
Hierarchical Part Segmentation. After segmentation, we can obtain the average feature of each part by finding the corresponding point-wise feature $\mathbf{f}_{p}$. We then directly employ hierarchical clustering on these average features to obtain hierarchical segmentation results. As shown in Figure 6, our hierarchical segmentation results effectively aggregate different parts at various levels, validating the effectiveness of our feature extraction method in accurately representing the information of each part. Compared to the results from PartField, our method’s aggregation better adheres to the relationships between parts.
Part Generation. Our results can also be applied to part generation as an instruction for splitting objects. Figure 6 demonstrates the exploded part generation results of HoloPartYang et al. (2025) when given the segmentation masks from SAMPart3D and ours. Our more accurate segmentation masks can significantly improve HoloPart’s performance, helping it generate cleaner and more precise parts.
Figure 6: The applications of our method including multi-prompts and hierarchical part segmentation and part generation.
4.4 Ablation Study
We first conduct ablation studies on our network architecture. The feature extractor can be any point cloud encoder, and we select the current state-of-the-art one that is Sonata. The IoU predictor is critical for selecting the best mask; without it, our method cannot function properly. We then focus on conducting ablation studies of our two-stage multi-head segmentor. As shown in Table 4, we evaluate four ablated models and our full method on the test set of our dataset. The first model contains only one segmentation head of the first stage. The second and third models respectively include only the first and second stages. The fourth model includes both stages but is trained without data augmentation. The last model is our full version. This progressive ablation study clearly demonstrates the importance of each component in our method. Notably, the difference between the second and third models is that the latter extracts a global feature during segmentation. The better performance metrics of the third model highlight the importance of this global feature. As shown in Figure 7, we also present the full segmentation results without using the NMS or flood fill algorithm in our automatic segmentation approach. The results highlight the necessity of these two steps, as without them, the segmentation masks are not complete, have unclear boundaries, and do not yield a reasonable number of parts. Figure 7 also visualizes the point-wise features of objects. The results show that for the same type of data, our method produces similar features for corresponding parts, while our features can capture more detailed geometry compared to PartField, such as the eyes and ears of the person on the right. This fully demonstrates the advantages of our method in extracting accurate features.
Figure 7: The ablation study of our automatic segmentation and the visualization of our point-wise feature.